How many distinguishable permutations with repeated elements are there for the number 20366650?


1
Expert's answer
2021-06-10T09:43:53-0400

The number of permutations of "n" elements taken "n" at a time, with "r_1" elements of one kind, "r_2" elements of another kind, and so on, is


"\dfrac{n!}{r_1!r_2!...r_k!}"

20366650

There are "n=8" digits. The digit repeats "r_1=2" times, the digit "6" repeats "r_2=3" times.

Then

"n=8, r_1=2, r_2=3"


"\dfrac{8!}{2!3!}=\dfrac{8(7)(6)(5)(4)}{1(2)(3)}=1120"